Given that x, x/2 and x/4 is a geometric progression. Find the sum of the 1st 11 terms of the GP.
Answer by greenestamps(13215)



Sum of the first 2 terms:
x + x/2 = x(1+1/2) = x(2-1/2)

Sum of the first 3 terms:
x+x/2+x/4 = x(1+1/2+1/4) = x(2-1/4)

Sum of the first 4 terms:
x+x/2+x/4+x/8 = x(1+1/2+1/4+1/8) = x(2-1/8)

...
look at the pattern...
...

Sum of the first 11 terms:
x+x/2+x/4+...+x/2^10 = ... = x(2-1/2^10) = x(2-1/1024) = (2047/1024)x